# A Heuristic Algorithm Based on Beam Search and Iterated Local Search for the Maritime Inventory Routing Problem

### v1

@@ Abstract
Maritime Inventory Routing Problem (MIRP) plays a crucial role in the integration of global maritime commerce levels. However, there are still no well-established methodologies capable of efficiently solving large MIRP instances or their variants due to the high complexity of the problem. The adoption of exact methods, typically based on Mixed Integer Programming (MIP), for daily operations is nearly impractical due to the CPU time required, as planning must be executed multiple times while ensuring high-quality results within acceptable time limits. Non-MIP-based heuristics are less frequently applied due to the highly constrained nature of the problem, which makes even the construction of an effective initial solution challenging. Papageorgiou et al. (2014) introduced a single-product MIRP as the foundation for MIRPLib, aiming to provide a collection of publicly available benchmark instances. However, only a few studies that propose new methodologies have been published since then. To encourage the use of MIRPLib and facilitate result comparisons, this study presents a heuristic approach that does not rely on mathematical optimization techniques to solve a deterministic, finite-horizon, single-product MIRP. The proposed heuristic combines a variation of a Beam Search algorithm with an Iterated Local Search procedure. Among the 72 instances tested, the developed methodology can improve the best-known solution for ten instances within an acceptable CPU time.
Keywords Inventory Routing Problem  $\cdot$ Hybrid Genetic Search  $\cdot$ Metaheuristics  $\cdot$ Vendor Managed Inventory
@@ 1 Introduction
Solving MIRPs is a challenging task due to its alignment with real-world complexities, such as variable demand rates, multi-product logistics, and the coordination of a central supplier with multiple consumers. Fagerholt et al. (2023) reviewed MIRP research from the past decade and proposed several future directions, including applications in energy transport, such as hydrogen, ammonia, LNG, and liquefied CO2, considering increasingly aggressive decarbonization goals. They also emphasize a key consideration when addressing this type of problem: the different planning levels, particularly between the tactical and operational levels. Strategic and tactical planning problems typically involve decisions with long-term impacts, while operational planning focuses on planning over shorter time horizons, ranging from a few days or weeks to a few months, and tends to involve much harder constraints. In addition, operational planning problems require tools capable of providing high-quality results in acceptable time frames. As a result, exact methods are generally not favored as they typically deliver optimal solutions only for small-scale instances of practical size. Another challenge in MIRP is the strong interdependence between routing decisions, where factors such as inventory levels and berth limitations trigger chain reactions throughout the supply chain.
@@
There is no single and well-defined definition of an MIRP, but rather numerous variants tailored to specific industrial applications (Papageorgiou et al. (2014), Andersson et al. (2010)). Given this diversity, standardizing the comparison of solution methods and establishing a taxonomy is essential. To address this issue, Papageorgiou et al. (2014) introduced the MIRPLib, a collection of publicly available benchmark single-product MIRP instances. A subset of these instances, called Group 2, serves as the focus of this study. These instances feature a planning horizon of over 120 periods, with no split pickups or deliveries. The problem setting includes a single production port, multiple consumer ports, and various classes of vessels.
In this work, our objective is to develop a pure heuristic approach for MIRP applied to instances of Group 2 by Papageorgiou et al. (2014) available in MIRPLib. The heuristic consists of a variation of the Beam Search (BS) algorithm with an Iterated Local Search (BS-ILS). Achieving this goal requires several challenging steps, including the development of a package to read MIRPLib instances and the creation of a novel solution representation.
The main contributions of this work are: (i) a comprehensive package designed to extend the applicability of MIRPLib; (ii) the introduction of an innovative solution representation for MIRP; and (iii) the development of a new approach using the BS-ILS algorithm to address a deterministic MIRP within reasonable CPU times.
This paper is organized as follows. Section 2 briefly reviews the literature on deterministic MIRP works, its variants, and heuristic applications are presented. Section 3 describes the deterministic MIRP considered in this paper. Section 4 presents the proposed BS-ILS heuristic. Section 5 reports and discusses the results of computational experiments. Finally, Section 6 concludes the work and provides directions for future research.
@@ 2 Literature Review
This section provides a concise definition of MIRP, along with an overview of key studies that use heuristic methods to address it.
2.1 Maritime Inventory Routing Problems
Ship routing plays a critical role in global maritime trade, integrating and optimizing supply chains to achieve lower costs and increased efficiency (Christiansen and Fagerholt, 2009). The MIRP combines ship routing and scheduling with inventory management. It involves determining the specific ports to be visited, the sequence of visits, and the quantities to load or unload at each port, while respecting inventory constraints and minimizing total costs (Sanghikian and Martinelli, 2020).
Real-world MIRP applications are often complex, incorporating variations such as supplier/producer or central consumer configurations, inventory constraints at specific ports, variable production/consumption rates, multiple products, and spot charters. These extensions increase computational effort, making the development of effective solution methods particularly challenging. Fagerholt et al. (2023) examined MIRP variants and extensions, addressing multiple products, finite planning horizons, speed optimization, LNG applications, uncertainty handling, and the integration of combined cargo and inventory routing.
Papageorgiou et al. (2014) introduced the MIRPLib to standardize MIRP definitions and support advanced formulations. They proposed a core model for single-product distribution, ensuring inventory constraints at all ports. Building on this foundation, additional MIRP variants were developed, including Group 2 instances — deterministic, long-horizon problems with one loading port and multiple discharge ports. This simplified model assumes that the vessels travel fully loaded to discharge regions and return empty. Detailed descriptions of the original Group 2 model are provided in Section 3.
@@ 2.2 Heuristic Approaches
Many authors have developed various methods to address this problem, including mathematical models, heuristics, metaheuristics, and matheuristics. Shaabani et al. (2023) and Fagerholt et al. (2023) provided a comprehensive review of the literature on MIRP and the algorithms used. They summarized and classified recent MIRP studies, emphasizing factors such as the size of the problem (e.g., the number of ports, vessels, products, and planning horizon) and the techniques applied to solve them. Out of the 34 studies listed by Shaabani et al. (2023) and the 53 studies listed by Fagerholt et al. (2023), 38 focus on MIRP solutions using heuristic techniques.
@@
Metaheuristic approaches have become important for exploring large search spaces. Christiansen et al. (2011) integrated a construction heuristic with a Genetic Algorithm (GA) for multi-product MIRPs. Siswanto et al. (2011a) proposed a multi-heuristic methodology that achieved global optimality for smaller instances. Expanding further, Siswanto et al. (2011b) developed a multi-heuristic strategy for a ship inventory routing and scheduling problem with non-dedicated compartments. More recently, Siswanto et al. (2019) proposed a GA based on multi-heuristics for an MIRP with multiple time windows. Sheikhtajian et al. (2020) studied uncertain LNG marine IRPs and proposed a metaheuristic that combines heuristics, GA, and fuzzy modeling to improve sensitivity analysis.
@@
In MIRPLib-specific studies, Papageorgiou et al. (2018) introduced several heuristics, such as RHH, $K$-opt heuristics, local branching, solution polishing and hybrid approaches, integrated with a matheuristic approach in mathematical programming to solve MIRPLib Group 2 instances. Their study resulted in new best-known solutions for 26 out of 70 instances that had not yet been proven optimal, and they established new best-known bounds for 56 instances. Munguía et al. (2019) presented the Parallel Alternating Criteria Search, an LNS-based heuristic that yielded strong results for Groups 1 and 2. Friske and Buriol (2020) implemented a multi-start heuristic with LNS for Group 1 instances, while Friske et al. (2022) employed a relax-and-fix algorithm followed by fix-and-optimize to refine MIRPLib solutions. Wang et al. (2023) introduced a bilevel parallel memetic algorithm for a maritime location inventory routing problem, tested in Group 1 instances, achieving near-best solutions with favorable deviations and faster computation times.
In contrast to the reviewed studies, which often rely on mathematical models, this study employs a heuristic approach that does not rely on mathematical optimization techniques. This method minimizes the dependency on complex decision models while maintaining practical solution quality, distinguishing itself through accessibility and applicability to real-world scenarios.
@@ 3 Problem Description
In a defined time horizon of a set $T$, a heterogeneous fleet of ships, classified into different vessel classes from set $VC$, each with distinct capacities $Q^{vc}$, is assigned to service the production ports of a set $J^{P}$ and the consumption ports of a set $J^{C}$. This operation must be done while ensuring that the minimum and maximum inventory limits in port $j\in J$ ($S_{j,t}^{min}$ and $S_{j,t}^{max}$, respectively) are respected. Each port $j\in J$ is characterized by an initial inventory $s_{j,0}$ and a specific production or consumption rate $d_{j,t}$ in period $t\in T$, along with an associated indicator parameter $\Delta_{j}$, which takes the value $+1$ for production ports and $-1$ for consumption ports. Berth limit ($B_{j}$) constraints also must be respected, which restrict the number of vessels that can attempt to load or discharge in a port $j\in J$ in a given period of time. In this MIRP, a vessel must travel at full capacity from a loading port to a discharging port, and empty from a discharging port to a loading port.
@@
The formulation follows an arc-flow mixed-integer linear programming model, where a set of nodes $\mathcal{N}$ represents the port-time pairs. A set of arcs $A^{vc}$ associated with each vessel class $vc\in VC$ defines the possible routes between nodes. Furthermore, a set of nodes $\mathcal{N}_{s,t}$ is defined, incorporating both the source $n_{s}$ and the sink $n_{t}$ nodes. The set of inter-regional arcs $A^{vc,ir}$ represents the arcs between the production and consumption regions. For each vessel class $vc\in VC$ and for each node $n\in\mathcal{N}_{s,t}$, the incoming and outgoing arcs are denoted by $RS^{vc}_{n}$ (for reverse star) and $FS^{vc}_{n}$ (for forward star), respectively, while the set of forward star inter-regional arcs is denoted by ($FS^{vc,ir}_{n}$).
@@
All routing costs are incorporated into $C^{vc}_{a}$ (cost for vessel of vessel class $vc\in VC$ to traverse the arc $A^{vc}$). Arc costs may vary depending on the type of arc (source, sink, waiting, inter-port). The distance $\delta_{j_{1},j_{2}}$ between ports $j_{1}$ and $j_{2}$, the cost per kilometer per vessel $C^{vc}_{km}$ according to its vessel class $vc\in VC$ and the fees $\pi_{j}$ of each port $j\in J$ are considered. Some discounts are also included: a reward for finishing the route before the time horizon ends $\rho$ and a discount $\gamma^{vc}$ for vessels of a vessel class $vc\in VC$ traveling empty from a consumption port to a production port. Inventory costs $P_{j,t}$ are included as penalties, accounting for the amount of production and consumption in each port $j\in J$ to respect inventory levels in period $t\in T$.
@@
The continuous decision variables $\alpha_{j,t}$ account for the amount to be produced or consumed in port $j\in J$ during period $t\in T$, and $s_{j,t}$ the quantity of inventory units in port $j\in J$ at the end of period $t\in T$. The integer decision variables $x_{a}^{vc}$ represent the number of vessels in a given vessel class $vc\in VC$ that traverse a specific arc $a\in A^{vc}$.
@@
$\sum\limits_{a\in FS_{n}^{vc,ir}}x_{a}^{vc}-\sum\limits_{a\in RS_{n}^{vc,ir}}x_{a}^{vc}=\begin{cases}1&\text{if }n=n_{s},\\ -1&\text{if }n=n_{t},\\ 0&\text{if }n\in\mathcal{N}\setminus\{n_{s},n_{t}\}\end{cases}$ $n\in\mathcal{N}_{s,t},\,vc\in VC$ (2)
@@
$x_{a}^{vc}\in\mathbb{Z_{+}}$ $v\in VC,\,a\in A^{vc}\setminus A^{vc,ir}.$ (8)
In this model, the objective function (3) minimizes the sum of transportation costs, and penalties for lost production and stockout. Constraints (2) ensure the flow balance of vessels within each vessel class. The minimum and maximum inventory limits in ports must be respected (Constraints (3) and (4)). Constraints (5) represent the berth limit constraints, while Constraints (6) ensure that at most one vessel per vessel class travels along a particular interregional arc. Constraints (7) and (8) define the variables’ domains.
@@ 4 Proposed Heuristic
The developed heuristic is a modified version of BS (Lowerre (1976), Ow and Morton (1988), Reddy (1977)), which serves as a constructive algorithm, complemented by local search procedures that improve all the generated solutions. Once these are concluded, an ILS (Lourenço et al., 2003) is applied to possibly refine the incumbent solution.
@@
Figure 1: Proposed Heuristic
@@ 4.1 Solution Representation
Solutions are represented by a single vector that stores port-vessel information, defined as a call, and the order of the vector’s calls corresponds to the chronological order of the ships’ visits to these ports. Thus, a ship requiring to service a call in a port must wait its turn in the unique list to perform the operation. Figure 2 shows an example of the solution representation, which is structured as a sequence of calls corresponding to the pair (port, vessel). The sequence is chronological, but only when it involves the same ports and vessels across the calls. For a call to occur, the preceding ones must already have taken place, with respect to the port and the vessel. In the given example, the first port serviced is port P1 by vessel V3. The second port to be serviced is also port P1 by vessel V1. This means that vessel V3 must end its operation before vessel V1 begins (considering that there is only one berth in port P1). The third call corresponds to port P3, served by vessel V2 and the fourth call corresponds to vessel V1 attending port P2. To ensure feasibility, a vessel must always travel between a production port and a consumption port, in either direction, so that there are no travels between ports of identical type.
Figure 2: Example of a solution representation Call(port, vessel).
Figure 3: Redundant solution under the single-list representation.
@@
It should be noted that our solution representation allows a given solution to have multiple representations, leading to a redundancy, as illustrated in Figure 3. Both represented solutions correspond to the same actual solution, as the calls involve distinct ships servicing different ports. Note that it is not straightforward to check which solutions are equivalent. A specific approach has been created to address this issue with some auxiliary structures, as follows. For each element in the solution vector, four auxiliary pointers are maintained, each storing the position of the last occurrence of the vessel, the last occurrence of the port, the next occurrence of the vessel, and the next occurrence of the port. This information is preprocessed and used during local search procedures to restrict them to non-redundant operations, avoiding unnecessary computation. In addition, the time for each vessel and port is also stored in order to speed up the solution evaluation.

### v2

@@ Abstract
Maritime Inventory Routing Problem (MIRP) plays a crucial role in the integration of global maritime commerce levels. However, there are still no well-established methodologies capable of efficiently solving large MIRP instances or their variants due to the high complexity of the problem. The adoption of exact methods, typically based on Mixed Integer Programming (MIP), for daily operations is nearly impractical due to the CPU time required, as planning must be executed multiple times while ensuring high-quality results within acceptable time limits. Non-MIP-based heuristics are less frequently applied due to the highly constrained nature of the problem, which makes even the construction of an effective initial solution challenging. Papageorgiou et al. (2014) introduced a single-product MIRP as the foundation for MIRPLib, aiming to provide a collection of publicly available benchmark instances. However, only a few studies that propose new methodologies have been published since then. To encourage the use of MIRPLib and facilitate result comparisons, this study presents a heuristic approach that does not rely on mathematical optimization techniques to solve a deterministic, finite-horizon, single-product MIRP. The proposed heuristic combines a variation of a Beam Search algorithm with an Iterated Local Search procedure. Among the 72 instances tested, the developed methodology can improve the best-known solution for 19 instances within an acceptable CPU time.
Keywords Maritime Inventory Routing Problem  $\cdot$ Heuristics  $\cdot$ Beam Search  $\cdot$ Iterated Local Search
@@ 1 Introduction
Solving MIRPs is a challenging task due to its alignment with real-world complexities, such as variable demand rates, multi-product logistics, and the coordination of a central supplier with multiple consumers. Fagerholt et al. (2023) reviewed MIRP research from the past decade and proposed several future directions, including applications in energy transport, such as hydrogen, ammonia, LNG, and liquefied CO2, considering increasingly aggressive decarbonization goals. They also emphasize a key consideration when addressing this type of problem: the different planning levels, particularly between the tactical and operational levels. Strategic and tactical planning problems typically involve decisions with long-term impacts, while operational planning focuses on planning over shorter time horizons, ranging from a few days or weeks to a few months, and tends to involve much harder constraints. In addition, operational planning problems require tools capable of providing high-quality results in acceptable time frames. As a result, exact methods are generally not favored as they typically deliver optimal solutions only for small-scale instances of practical size. Another challenge involving MIRP is the strong interdependence between routing decisions, where factors such as inventory levels and berth limitations trigger chain reactions throughout the supply chain.
@@
There is no single and well-defined definition of an MIRP, but rather numerous variants tailored to specific industrial applications (Papageorgiou et al. (2014), Andersson et al. (2010)). Given this diversity, standardizing the comparison of solution methods and establishing a taxonomy is essential. To address this issue, Papageorgiou et al. (2014) introduced the MIRPLib, a collection of publicly available single-product MIRP benchmark instances. A subset of these instances, called Group 2, serves as the focus of this study. These instances feature a planning horizon of over 120 periods, with no split pickups or deliveries. The problem setting includes a single production port, multiple consumer ports, and various classes of vessels.
In this work, our objective is to develop a pure heuristic approach for MIRP applied to instances of Group 2 by Papageorgiou et al. (2014) available in MIRPLib. The heuristic consists of a variation of the Beam Search (BS) algorithm (Lowerre (1976), Ow and Morton (1988), Reddy (1977)) with an Iterated Local Search (ILS) (Lourenço et al., 2003) (BS-ILS).
The main contributions of this work are: (i) the introduction of an innovative solution representation for MIRP; (ii) the development of a new approach using the BS-ILS algorithm to address a deterministic MIRP within reasonable CPU times; and (iii) a Julia package designed to support the applicability of the MIRPLib.
This paper is organized as follows. Section 2 briefly reviews the literature on deterministic MIRP works and its variants, and presents heuristic applications. Section 3 describes the deterministic MIRP considered in this paper. Section 4 presents the proposed BS-ILS heuristic. Section 5 reports and discusses the results of computational experiments. Finally, Section 6 concludes the work and provides directions for future research.
@@ 2 Literature Review
2.1 Maritime Inventory Routing Problem
The MIRP refers to the strategic coordination of maritime transportation and inventory management, integrating logistics decisions with dynamic inventory levels at both origins and destinations. It requires planning vessel routes, determining port call sequences and deciding product quantities for loading and unloading, while ensuring that inventory limits are respected (Sanghikian and Martinelli, 2020).
Real-world MIRP applications are often complex, incorporating variations such as supplier/producer or central consumer configurations, inventory constraints at specific ports, variable production/consumption rates, multiple products, and spot charters. To support the development and comparison of solution methods for these variants, standardized instances and modeling frameworks are essential. In this context, Papageorgiou et al. (2014) introduced the MIRPLib to standardize MIRP definitions and support advanced formulations. They proposed a core model for single-product distribution, ensuring inventory constraints at all ports. Building on this foundation, additional MIRP variants were developed, including Group 2 instances — deterministic, long-horizon problems with one loading port and multiple discharge ports. This simplified model assumes that the vessels travel fully loaded to discharge regions and return empty. Detailed descriptions of the original Group 2 model are provided in Section 3.
@@ 2.2 Heuristic Approaches
Many authors have developed methods to address MIRP, including mathematical models, heuristics, metaheuristics, and matheuristics. Shaabani et al. (2023) and Fagerholt et al. (2023) provided a comprehensive review of the literature on MIRP and the algorithms used. They summarized and classified recent MIRP studies, emphasizing factors such as the size of the problem (e.g., the number of ports, vessels, products, and planning horizon) and the techniques applied to solve them. Out of the 34 studies listed by Shaabani et al. (2023) and the 53 studies listed by Fagerholt et al. (2023), 38 focus on MIRP solutions using heuristic techniques.
@@
Metaheuristic approaches have become important for exploring large search spaces. Christiansen et al. (2011) integrated a construction heuristic with a Genetic Algorithm (GA) for multi-product MIRPs. Siswanto et al. (2011a) proposed a multi-heuristic methodology that achieved optimality for smaller instances. Expanding further, Siswanto et al. (2011b) developed a multi-heuristic strategy for a ship inventory routing and scheduling problem with non-dedicated compartments. More recently, Siswanto et al. (2019) proposed a GA based on multi-heuristics for an MIRP with multiple time windows. Sheikhtajian et al. (2020) studied uncertain LNG marine IRPs and proposed a metaheuristic that combines heuristics, GA, and fuzzy modeling to improve sensitivity analysis.
@@
In MIRPLib-specific studies, Papageorgiou et al. (2018) introduced several heuristics, such as RHH, $K$-opt heuristics, local branching, solution polishing and hybrid approaches, that combine metaheuristics with mathematical programming to solve MIRPLib’s Group 2 instances. Their study resulted in new best-known solutions for 26 out of 70 instances that had not yet been proven optimal, and they achieved new best-known lower bounds for 56 instances. Munguía et al. (2019) presented the Parallel Alternating Criteria Search, an LNS-based heuristic that yielded strong results for Groups 1 and 2.
Friske and Buriol (2020) implemented a multi-start heuristic with LNS for Group 1 instances, while Friske et al. (2022) employed a relax-and-fix algorithm followed by fix-and-optimize to refine MIRPLib’s Group 1 solutions. Wang et al. (2023) introduced a bilevel parallel memetic algorithm for a maritime location inventory routing problem, tested in Group 1 instances, achieving near-best solutions with favorable deviations and faster computation times.
In contrast to the reviewed studies, which often rely on mathematical models, this study employs a heuristic approach that avoids mathematical programming techniques. This method minimizes the dependency on complex decision models while maintaining practical solution quality, distinguishing itself through accessibility and applicability to real-world scenarios.
@@ 3 Problem Description
In a defined time horizon $T$, a heterogeneous fleet of ships, classified into different vessel classes $VC$, each with distinct capacities $Q^{vc}$, is assigned to service the production ports $J^{P}$ and the consumption ports $J^{C}$. This operation must be done while ensuring that the minimum and maximum inventory limits in port $j\in J$ ($S_{j,t}^{min}$ and $S_{j,t}^{max}$, respectively) are respected. Each port $j\in J$ is characterized by an initial inventory $s_{j,0}$ and a specific production or consumption rate $d_{j,t}$ in period $t\in T$, along with an associated indicator parameter $\Delta_{j}$, which takes the value $1$ for production ports and $-1$ for consumption ports. Berth limit ($B_{j}$) constraints must also be respected, restricting the number of vessels that can attempt to load or discharge in a port $j\in J$ in a given period of time. In this MIRP, a vessel must travel at full capacity from a loading port to a discharging port, and empty from a discharging port to a loading port.
@@
The formulation follows an arc-flow mixed-integer linear programming model, where a set of nodes $\mathcal{N}$ represents the port-time pairs. A set of arcs $A^{vc}$ associated with each vessel class $vc\in VC$ defines the possible routes between nodes. Furthermore, a set of nodes $\mathcal{N}_{s,t}$ is defined, including both the source $n_{s}$ and the sink $n_{t}$ nodes. The set of inter-regional arcs $A^{vc,ir}$ represents the arcs between the production and consumption regions. For each vessel class $vc\in VC$ and for each node $n\in\mathcal{N}_{s,t}$, the incoming and outgoing arcs are denoted by $RS^{vc}_{n}$ (for reverse star) and $FS^{vc}_{n}$ (for forward star), respectively, while the set of forward star inter-regional arcs is denoted by ($FS^{vc,ir}_{n}$).
@@
All routing costs are incorporated into $C^{vc}_{a}$ (cost for vessel of vessel class $vc\in VC$ to traverse the arc $a\in A^{vc}$). Arc costs may vary depending on the type of arc (source, sink, waiting, inter-port). The distance $\delta_{j_{1},j_{2}}$ between ports $j_{1}$ and $j_{2}$, the cost per kilometer per vessel $C^{vc}_{km}$ according to its vessel class $vc\in VC$ and the fees $\pi_{j}$ of each port $j\in J$ are considered. Some discounts are also included, a reward for finishing the route before the time horizon ends $\rho$ and a discount $\gamma^{vc}$ for vessels of a vessel class $vc\in VC$ traveling empty from a consumption port to a production port. Inventory costs $P_{j,t}$ are included as penalties, reflecting production and consumption in each port $j\in J$ to comply with inventory constraints in each period $t\in T$.
@@
Integer decision variables $x_{a}^{vc}$ represent the number of vessels in a given vessel class $vc\in VC$ that traverse a specific arc $a\in A^{vc}$. The continuous decision variables $\alpha_{j,t}$, representing a spot charter, represent the amount to be produced or consumed in port $j\in J$ during period $t\in T$, and $s_{j,t}$ the quantity of inventory units in port $j\in J$ at the end of period $t\in T$.
@@
$\sum\limits_{a\in FS_{n}^{vc}}x_{a}^{vc}-\sum\limits_{a\in RS_{n}^{vc}}x_{a}^{vc}=\begin{cases}1&\text{if }n=n_{s},\\ -1&\text{if }n=n_{t},\\ 0&\text{if }n\in\mathcal{N}\setminus\{n_{s},n_{t}\}\end{cases}$ $n\in\mathcal{N}_{s,t},\,vc\in VC$ (2)
@@
$x_{a}^{vc}\in\{0,1\}$ $vc\in VC,\,a\in A^{vc,ir}$ (7)
$x_{a}^{vc}\in\mathbb{Z_{+}}$ $vc\in VC,\,a\in A^{vc}\setminus A^{vc,ir}.$ (8)
In this model, the objective function (3) minimizes the sum of transportation costs, and penalties for lost production and stockout. Constraints (2) ensure the flow balance of vessels within each vessel class. The minimum and maximum inventory limits at ports must be respected, while also ensuring the continuity of inventory flow, as defined by Constraints (3) and (4), respectively. Constraints (5) represent the berth limit constraints. Constraints (6) ensure that the spot charter variable is always non-negative. Constraints (7) ensure that at most one vessel per vessel class travels along a particular interregional arc. Finally, Constraints (8) define the variables’ domains for the remaining arcs.
Solving MIRP using exact mathematical models is often impractical due to their high computational cost. As the problem size increases, particularly in real-world instances, solution times become unmanageable, limiting the applicability of exact approaches. Consequently, exact methods are limited to small or simplified cases, highlighting the need for scalable alternatives like heuristics or metaheuristics that can provide high-quality solutions within a reasonable CPU time.
@@ 4 Proposed Heuristic
The developed heuristic is a modified version of Beam Search (Lowerre (1976), Ow and Morton (1988), Reddy (1977)), which serves as a constructive algorithm, complemented by local search procedures that improve all the generated solutions. Once these are concluded, an Iterated Local Search (Lourenço et al., 2003) is applied to possibly refine the incumbent solution.
@@
Figure 1: Proposed Heuristic.
@@ 4.1 Solution Representation
The solutions are represented by a single vector that stores port-vessel information, defined as a call, and the order of the calls in the vector corresponds to the chronological order of the ships’ visits to these ports. For a call to occur, the preceding ones must already have taken place, with respect to the port and the vessel. Thus, a ship requiring to service a call in a port must wait its turn in the list to perform the operation. The sequence is chronological, but only when it involves the same ports or vessels across the calls. For example, if the first port to be serviced is port P1 by vessel V3, followed by port P1 again, but this time by vessel V1, vessel V3 must end its operations at port P1 before vessel V1 can begin (considering that there is only one berth available at port P1). On the other hand, if the next call is unrelated, it may occur prior to the last one. To ensure feasibility, vessels must always travel between a production port and a consumption port, in either direction. This rule prevents vessels from traveling between ports of the same type, ensuring operational consistency.
This solution representation allows a given solution to have multiple representations, leading to redundancy. Two solutions can correspond to the same actual solution, as the calls involve distinct ships servicing different ports. For example, a solution may evaluate (P1,V1) followed by (P3,V2), while another solution evaluates (P3,V2) first and then (P1,V1). Despite the difference in visit order, both solutions correspond to the same operation in real life. This illustrates that permutations may not affect the final outcome and can be considered equivalent in the problem context.
@@
Note that it is not straightforward to check which solutions are equivalent. A specific approach has been created to address this issue with some auxiliary structures as follows. For each element in the vector, four auxiliary pointers are maintained, each storing the position of the last occurrence of the vessel, the last occurrence of the port, the next occurrence of the vessel, and the next occurrence of the port. This information is preprocessed and used during local search procedures to restrict their moves to non-redundant operations, avoiding unnecessary computation. In addition, the time for each vessel and port is also preprocessed to speed up the solution evaluation.
4.2 Solution Evaluation
With the aim of evaluating a solution, each call keeps key information such as the most recent service time at the respective port, the last service time of the vessel, inventory levels, the time of the next violation and the accumulated solution costs (including both routing and inventory components), in addition to the previously described pointers. Depending on the type of change introduced, whether by local search, ILS perturbation, or the greedy algorithm itself, any change in routing leads to the evaluation, responsible for recalculating the relevant solution information.
Upon evaluation of a call, the arrival and operation times of the vessel in the port are calculated considering berth constraints and avoiding potential infeasibility. If the operation occurs within acceptable limits and no constraint is violated, then no penalty is incurred, and inventory levels at the ports are updated accordingly. However, if a violation occurs prior to the scheduled operation, penalty costs are applied, and inventory levels are corrected to the admissible limit (upper or lower), depending on the port type (production or consumption).
Subsequently, routing costs are calculated and the planning horizon is checked. Calls exceeding the horizon are discarded, and the final inventory violations and total routing costs are recalculated. At the end of each evaluation, all relevant information is updated to ensure consistency, allowing for the next evaluation in upcoming movements.